OpenStudy (girlgoyleh.):

How would i figure this out? 5(x-1) = x+11

KjSaif (kjsaif):

Simplify both sides of the equation.

KjSaif (kjsaif):

(5)(x)+(5)(−1)=x+11(Distribute) 5x + −5 = x+11 5x − 5 = x+ 11

KjSaif (kjsaif):

Subtract x from both sides.

KjSaif (kjsaif):

5x − 5 − x = x+11−x 4x − 5=11

KjSaif (kjsaif):

Add 5 to both sides.

KjSaif (kjsaif):

4x − 5 + 5=11+5 4x=16

KjSaif (kjsaif):

Divide both sides by 4. :)

KjSaif (kjsaif):

4x/4 = 16/4

KjSaif (kjsaif):

x=4

KjSaif (kjsaif):

Hope this helps! @girlgoyleh.

OpenStudy (girlgoyleh.):

It did a lot, thank you very much :)

KjSaif (kjsaif):

:)